# Investigating symptom duration using current status data: a case study of post-acute COVID-19 syndrome

### v2

@@ Abstract
For infectious diseases, characterizing symptom duration is of clinical and public health importance. Symptom duration may be assessed by surveying infected individuals and querying symptom status at the time of survey response. For example, in a SARS-CoV-2 testing program at the University of Washington, participants were surveyed at least 28 days after testing positive and asked to report current symptom status. This study design yielded current status data: outcome measurements for each respondent consisted only of the time of survey response and a binary indicator of whether symptoms had resolved by that time. Such study design benefits from limited risk of recall bias, but analyzing the resulting data necessitates tailored statistical tools. Here, we review methods for current status data and describe a novel application of modern nonparametric techniques to this setting. The proposed approach is valid under weaker assumptions compared to existing methods, allows use of flexible machine learning tools, and handles potential survey nonresponse. From the university study, we estimate that 19% of participants experienced ongoing symptoms 30 days after testing positive, decreasing to 7% at 90 days. Female sex, history of seasonal allergies, fatigue during acute infection, and higher viral load were associated with slower symptom resolution.
@@ Introduction
The persistence of COVID-19 symptoms is a feature of the natural history of SARS-CoV-2 infection. Early in the COVID-19 pandemic, it became clear that many infected individuals suffered symptoms continuing beyond the acute phase of the disease [1, 2, 3]. Post-acute sequelae are not unique to SARS-CoV-2; many viruses, including the Ebola, Epstein-Barr, chikungunya, dengue, polio, and SARS viruses, among others, have been linked to post-acute infection syndromes [4]. There is now a substantial body of literature on the post-acute sequelae of COVID-19 (PASC), colloquially known as long COVID, but there is no universally agreed upon case definition for PASC. Recently, the National Academies proposed defining PASC as an infection-associated chronic condition persisting for at least three months after acute SARS-CoV-2 infection [5]. The World Health Organization instead suggests that a patient can be considered to experience PASC if, three months after infection, they have experienced symptoms lasting at least two months that are not explainable by an alternative diagnosis [6].
@@
Since there is no consensus on the duration of symptoms warranting a PASC diagnosis, it may be informative to study the distribution of time from SARS-CoV-2 infection — or, more feasibly, a suitable proxy such as onset of symptoms or confirmation of infection by testing — until symptom resolution. In this article, we are motivated by the problem of studying the duration of COVID-19 symptoms using data collected as part of the Husky Coronavirus Testing (HCT) initiative. From September $2020$ to June $2023$, the HCT study enrolled 40,000 participants at the University of Washington (UW) in Washington state [7]. Beginning in December $2021$, roughly at the beginning of the initial wave of Omicron variant infections, participants who reported a SARS-CoV-2 infection confirmed by either polymerase chain reaction (PCR) test or at-home rapid antigen test were invited to complete a follow-up survey regarding the persistence of symptoms in the months following infection. The aim of this survey was to estimate how common persistent post-acute COVID-19 symptoms are, and to identify risk factors for persistence in a population of university affiliates. Notably, this population is predominantly young and highly vaccinated, with relatively few preexisting comorbidities. In the survey, participants were asked whether, at the time of survey completion, they continued to experience symptoms associated with their recent infection.
@@
The exact time from infection until symptom resolution was not observed for any participant in the HCT long COVID study. Instead, for each survey respondent, in addition to baseline covariate information collected at or before the test date, the available data consist of the time elapsed from positive SARS-CoV-2 test until survey response — referred to as the survey response time — and a binary variable indicating whether the respondent’s symptoms had resolved as of the survey response time — referred to as the event indicator. Each participant was observed only at a single survey response time. Response times in this study ranged from $28$ days (the earliest time after a positive test at which surveys were sent) to $286$ days. Because by design investigators recorded only the current status of a participant’s symptom persistence at their response time, this sampling scheme, which represents an extreme form of interval censoring, is referred to as current status sampling.
@@
Current status data arise and have been analyzed in a number of epidemiological settings, including partner studies of HIV transmission [8], cross-sectional studies for estimating the age of onset of a non-fatal disease [9, 10], and demographic studies of age at weaning [11, 12, 13, 14, 15]. Designs analogous to the HCT long COVID study, in which participants are surveyed following a health-related event and queried regarding the presence of sequelae, have been used, for example, to study outcomes following radiation therapy [16], joint surgery [17, 18], diagnosis with light chain amyloidosis [19], onset of schizophrenia [20], and infection with Ebola virus [21]. In fact, previous studies of persistent COVID-19 symptoms have also used a similar design [22, 23, 24, 25]. Such designs may be appealing because they are simple to implement, economical, and most importantly, less susceptible to recall bias than retrospective designs in which participants are asked to recall symptom status at a specific date in the past [26]. Nevertheless, the presence of current status data may be easily missed; many of the aforementioned studies treated occurrence of the outcome as a binary variable, ignoring variation in response times. Recognizing the presence of current status data is vital, since their analysis requires specialized statistical techniques [27].
@@
Traditional analysis techniques for current status data require that survey response times be uninformative of the symptom resolution process. In this paper, we show that recent methodological advances we have made in causal inference under shape constraints can be adapted to make statistical inference on the time-to-symptom-resolution distribution allowing some response time informativeness. The proposed approach allows us to use flexible machine learning tools and thereby minimize the need for restrictive modeling assumptions. Furthermore, our method accounts for potential nonresponse, which is ubiquitous in survey studies such as the HCT long COVID study, where over half of survey recipients had not responded when the study concluded. To our knowledge, the link between the methodology we build upon here, termed causal isotonic regression (CIR), and the analysis of current status data has not been made before; we aim to highlight the utility of this approach. Using our method, we provide substantive results regarding persistent COVID-19 symptoms in the HCT population during the early Omicron period of the SARS-CoV-2 pandemic. We also investigate baseline patient characteristics associated with time to resolution, that is, risk factors for delayed symptom resolution, using existing regression methods for current status data.
@@ Methods
In an ideal study with unlimited follow-up, we would observe for each participant the survey response time $Y_{*}$ — the time elapsed between positive SARS-CoV-2 test and survey response — and the event indicator $\Delta_{*}$ — a participant’s symptom status at the time of survey response — with $\Delta_{*}=1$ representing symptom resolution by time $Y_{*}$ and $\Delta_{*}=0$ otherwise. We denote by $c_{0}$ the investigator-determined time (since the positive test) at which follow-up ends for a participant, and consider as a nonrespondent any participant without a response by time $c_{0}$. We set $Y=\min(Y_{*},c_{0})$ as the smaller of the survey response time and $c_{0}$, and denote by $\Delta=\mathbbm{1}(Y_{*}<c_{0})\Delta_{*}$ the observed response indicator, which equals $1$ if a participant responded and reported resolution and 00 otherwise. The available data are considered to be $n$ independent and identically distributed observations $(W_{1},Y_{1},\Delta_{1}),(W_{2},Y_{2},\Delta_{2}),\ldots,(W_{n},Y_{n},\Delta_{n})$.
@@
Traditionally, under current status sampling, the survival function has been estimated based on the nonparametric maximum likelihood estimator (NPMLE), which can be computed using the pool adjacent violators algorithm used for isotonic (i.e., monotone) regression [28]. The key idea behind this approach is that estimation of the survival function can be reframed as a regression problem under a monotonicity constraint: the probability of symptom resolution by time $t$ is simply the proportion of individuals who have resolved symptoms among those with survey response time $t$, and this probability is necessarily nondecreasing as a function of survey response time. This estimator has the advantage of being distribution-free, in the sense that its validity does not rely on any assumption about the underlying shape of the true survival function, and of enabling the construction of valid confidence intervals derived from its large-sample properties [29]. Nevertheless, an important limitation of the NPMLE approach is that it requires the survey response process to be entirely uninformative, in the sense that the response time itself does not carry information about the event time. This condition may fail to hold in practical settings, as there may exist baseline covariates — analogous to confounders in a causal inference setting — that inform the relationship between the event indicator and the survey response time. For example, participants with substantial comorbidities may be more likely to respond promptly to the survey and to have greater symptom persistence. We argue here that the framework of CIR introduced in Westling et al. [30], which was developed to estimate a monotone causal dose-response curve, can be directly used to estimate the survival function with current status data under weaker conditions. Specifically, as opposed to the traditional NPMLE approach, this method only requires independence between the survey response and event times within strata defined by the recorded baseline covariates. Such independence condition is typically much more realistic, and thereby makes the analysis less likely to yield misleading conclusions due to assumption violations.
@@
are then obtained and regressed against $Y_{1},Y_{2},\ldots,Y_{n}$ over the domain $[t_{0},t_{1}]$ using isotonic regression (see Supplementary Material for technical details). We denote the resulting estimator of $\psi(t)$ as $\psi_{n}(t)$. Use of these pseudo-outcomes rather than the simple event indicators accounts for possible dependence between $\Delta_{*}(t)$ and $Y_{*}$, as long as this dependence is explained entirely by $W$. Roughly speaking, if at least one of $\mu$ and $g$ is estimated consistently, then $\psi_{n}(t)$ is itself consistent for $\psi(t)$, and if both $\mu$ and $g$ are estimated consistently, $n^{1/3}\{\psi_{n}(t)-\psi(t)\}$ tends in distribution to a random variable with a scaled Chernoff distribution [31]. This well-characterized large-sample distribution allows for construction of confidence intervals with correct coverage. A key strength of the CIR approach is that it allows for flexible estimation of $\mu$ and $g$ — using machine learning, for example — without sacrificing statistical inference. This reduces the risk of systematic bias due to the use of misspecified parametric models.
@@
We note that performing the CIR procedure using data from survey respondents alone — in other words, performing a complete case analysis — could result in bias, since whether a participant is considered a respondent is determined by their value of $Y_{*}$, which may depend on $W$ and thus be informative for $\Delta_{*}$. Symbolically, the estimand in a complete case analysis would be $1-E\{E(\Delta\,|\,W,Y=y)\,|\,Y<c_{0}\}$ rather than $1-E\{E(\Delta\,|\,W,Y=y)\}$. Therefore, fully removing nonrespondents — that is, participants with $Y_{*}\geq c_{0}$ — may induce selection bias, as the distribution of $W$ may differ among survey respondents and nonrespondents. To emphasize the difference between the complete case CIR and our proposed method, which includes all survey recipients, we refer to the latter as extended CIR.
@@
It is also of interest to understand the association between baseline covariates $W$ and symptom duration $T$ in order to identify potential predictors of delayed symptom resolution. In the Supplementary Material, we review regression models that have been adapted for use with current status data and discuss the difficulties that arise in making inference using these models. Focusing on the Cox model, we conduct a simulation study to evaluate the performance of the nonparametric bootstrap for making inference, finding that Wald-type bootstrap confidence intervals based on a normal approximation demonstrate good performance in modest sample sizes. The Cox model requires no special adaptation to account for nonresponse, as long as the regression model is correctly specified and includes all covariates that inform the response mechanism [32].
@@
The HCT study was conducted at UW, a large public university in Washington state with a population composed of approximately 60,000 students and 32,000 faculty and staff. Beginning in autumn 2020, the HCT program provided free, voluntary PCR testing for SARS-CoV-2 and symptom surveillance for university affiliates. Individuals were sent daily messages via text or email to ask about new symptoms, exposures, and high-risk behaviors, any of which would trigger an invitation for PCR testing. Study participants were also able to walk-in for testing at any time, and for any reason, without an invitation. Once antigen testing became widely available in March $2022$, individuals were asked on their daily messages to report any antigen test results. The inclusion criteria for HCT eligibility are described in the Supplementary Material. Informed consent was signed electronically at the time of enrollment.
For this analysis, we used data collected from December $2021$ to June $2023$. During this time period, participants who had a positive or inconclusive PCR test or who reported a positive rapid antigen test for SARS-CoV-2 were invited to participate in a follow-up survey to assess for ongoing symptoms of COVID-19 a minimum of four weeks after the positive test.
@@
    Long COVID survey: Starting in December $2021$, all participants who tested positive or inconclusive for SARS-CoV-2 by PCR testing or those who attested to having had a positive rapid antigen test were contacted a minimum of $28$ days after their first positive test date and invited to participate in the long COVID survey. Each participant was allowed to fill out the survey only once per academic year. For participants with multiple responses, only the first is included in this analysis. The survey included the question “Are you still experiencing symptoms related to your COVID-19 illness (e.g., brain fog, fatigue, headaches, etc.)?” Each participant was free to choose when to complete the survey. Data collection was terminated on June $23$, $2023$.
@@
We implemented the extended CIR method using flexible machine learning tools to estimate $\mu$ and $g$. To estimate $\mu$, we used the Super Learner [33, 34], a form of stacked regression, with the optimal linear combination of learners selected to minimize $10$-fold cross-validated mean squared error — see Supplementary Material for the algorithm library. To estimate $\pi$, we used the haldensify software package [35] with $10$-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve, which is required for confidence interval construction, was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. Confidence intervals with nominal $95$% pointwise coverage were constructed using the quantiles of the Chernoff distribution.
@@
We aimed to estimate the survival function for symptom resolution between $t_{0}=30$ and $t_{1}=90$ days after the positive test. We therefore chose the maximum follow-up time to be $c_{0}=120$; participants with positive tests after February $24$, $2023$ were excluded to ensure all individuals in the analysis cohort had at least $120$ days of follow-up. On one hand, large values of $c_{0}$ result in more individuals excluded from the study period but make $t_{1}$ further away from the largest possible observed survey response times, where the statistical performance of isotonic regression and related procedures is known to degrade. On the other hand, small values of $c_{0}$ result in fewer individuals excluded from the study period but make $t_{1}$ closer to the largest possible observed survey response times. Our choice of $c_{0}$ was meant to achieve a balance between these competing considerations, and resulted in $t_{1}$ lying at approximately the $97.5$th percentile of observed survey response times; such choice performed well in a simulation study — see Supplementary Material for details.
@@
To investigate the association between baseline risk factors and time from infection to symptom resolution, we fit a Cox proportional hazards regression model accounting for current status sampling using the icenReg software package [36]. Main terms were included in the model for each baseline covariate described above. Standard error estimates were calculated using a nonparametric bootstrap with $1000$ bootstrap replicates, and used to construct $95$% Wald-type confidence intervals and perform two-sided tests of the null regression coefficient hypothesis at a $0.05$ level. For categorical risk factors, multivariate Wald tests were used to test the joint null hypothesis that all of the corresponding regression coefficients were zero.

### v3

@@ Abstract
For infectious diseases, characterizing symptom duration is of clinical and public health importance. Symptom duration may be assessed by surveying infected individuals and querying symptom status at the time of survey response. For example, in a SARS-CoV-2 testing program at the University of Washington, participants were surveyed at least $28$ days after testing positive and asked to report current symptom status. This study design yielded current status data: outcome measurements for each respondent consisted only of the time of survey response and a binary indicator of whether symptoms had resolved by that time. Such study design benefits from limited risk of recall bias, but analyzing the resulting data necessitates tailored statistical tools. Here, we review methods for current status data and describe a novel application of modern nonparametric techniques to this setting. The proposed approach is valid under weaker assumptions compared to existing methods, allows use of flexible machine learning tools, and handles potential survey nonresponse. From the university study, under an assumption that the survey response time is conditionally independent of symptom resolution time within strata of measured covariates, we estimate that $19$% of participants experienced ongoing symptoms $30$ days after testing positive, decreasing to $7$% at $90$ days. We assess the sensitivity of these results to deviations from conditional independence, finding the estimates to be more sensitive to assumption violations at 30 days compared to 90 days. Female sex, fatigue during acute infection, and higher viral load were associated with slower symptom resolution.
@@ Introduction
The persistence of COVID-19 symptoms is a feature of the natural history of SARS-CoV-2 infection. Early in the COVID-19 pandemic, it became clear that many infected individuals suffered symptoms continuing beyond the acute phase of the disease (Goërtz et al.,, 2020; Liang et al.,, 2020; van den Borst et al.,, 2021). Post-acute sequelae are not unique to SARS-CoV-2; many viruses, including the Ebola, Epstein-Barr, chikungunya, dengue, polio, and SARS viruses, among others, have been linked to post-acute infection syndromes (Choutka et al.,, 2022). There is now a substantial body of literature on the post-acute sequelae of COVID-19 (PASC), colloquially known as long COVID, but there is no universally agreed upon case definition for PASC. Recently, the National Academies proposed defining PASC as an infection-associated chronic condition persisting for at least three months after acute SARS-CoV-2 infection (National Academics of Science, Engineering, and Medicine,, 2024). The World Health Organization instead suggests that a patient can be considered to experience PASC if, three months after infection, they have experienced symptoms lasting at least two months that are not explainable by an alternative diagnosis (Soriano et al.,, 2022).
@@
Since there is no consensus on the duration of symptoms warranting a PASC diagnosis, it may be informative to study the distribution of time from SARS-CoV-2 infection — or, more feasibly, a suitable proxy such as onset of symptoms or confirmation of infection by testing — until symptom resolution. In this article, we are motivated by the problem of studying the duration of COVID-19 symptoms using data collected as part of the Husky Coronavirus Testing (HCT) initiative. From September $2020$ to June $2023$, the HCT study enrolled 40,000 participants at the University of Washington (UW) in Washington state (Weil et al.,, 2021). Beginning in December $2021$, roughly at the beginning of the initial wave of Omicron variant infections, participants who reported a SARS-CoV-2 infection confirmed by either polymerase chain reaction (PCR) test or at-home rapid antigen test were invited to complete a follow-up survey regarding the persistence of symptoms in the months following infection. The aim of this survey was to estimate how common persistent post-acute COVID-19 symptoms are, and to identify risk factors for persistence in a population of university affiliates. Notably, this population is predominantly young and highly vaccinated, with relatively few preexisting comorbidities. In the survey, participants were asked whether, at the time of survey completion, they continued to experience symptoms associated with their recent infection.
@@
The exact time from infection until symptom resolution was not observed for any participant in the HCT long COVID study. Instead, for each survey respondent, in addition to baseline covariate information collected at or before the test date, the available data consist of the time elapsed from positive SARS-CoV-2 test until survey response — referred to as the survey response time — and a binary variable indicating whether the respondent’s symptoms had resolved as of the survey response time — referred to as the event indicator. Each participant was observed only at a single survey response time. Response times in this study ranged from $28$ days (the earliest time after a positive test at which surveys were sent) to $257$ days. Because by design investigators recorded only the current status of a participant’s symptom persistence at their response time, this sampling scheme, which represents an extreme form of interval censoring, is referred to as current status sampling.
@@
Current status data arise and have been analyzed in a number of epidemiological settings, including partner studies of HIV transmission (Jewell and Shiboski,, 1990), cross-sectional studies for estimating the age of onset of a non-fatal disease (Keiding,, 1991; Keiding et al.,, 1996), and demographic studies of age at weaning (Grummer-Strawn,, 1993; Lutter et al.,, 2011; Rinaldi et al.,, 2019; Khan et al.,, 2020; Harvey et al.,, 2022). Designs analogous to the HCT long COVID study, in which participants are surveyed following a health-related event and queried regarding the presence of sequelae, have been used, for example, to study outcomes following radiation therapy (Head et al.,, 2017), joint surgery (Pinsker et al.,, 2016; Lim et al.,, 2022), diagnosis with light chain amyloidosis (D’Souza et al.,, 2023), onset of schizophrenia (Fan et al.,, 2021), and infection with Ebola virus (Mattia et al.,, 2016). In fact, previous studies of persistent COVID-19 symptoms have also used a similar design (Ganesh et al.,, 2021; Robineau et al.,, 2022; Imoto et al.,, 2022; Phen et al.,, 2024). Such designs may be appealing because they are simple to implement, economical, and most importantly, less susceptible to recall bias than retrospective designs in which participants are asked to recall symptom status at a specific date in the past (Landry et al.,, 2023). Nevertheless, the presence of current status data may be easily missed; many of the aforementioned studies treated occurrence of the outcome as a binary variable, ignoring variation in response times. Recognizing the presence of current status data is vital, since their analysis requires specialized statistical techniques (Jewell and van der Laan,, 2003).
@@
Traditional analysis techniques for current status data require that survey response times be uninformative of the symptom resolution process. In this paper, we show that recent methodological advances we have made in causal inference under shape constraints can be adapted to make statistical inference on the time-to-symptom-resolution distribution allowing some response time informativeness. The proposed approach allows us to use flexible machine learning tools and thereby minimize the need for restrictive modeling assumptions. Furthermore, our method accounts for potential nonresponse, which is ubiquitous in survey studies such as the HCT long COVID study, where over half of survey recipients had not responded when the study concluded. To our knowledge, the link between the methodology we build upon here, termed causal isotonic regression (CIR), and the analysis of current status data has not been made before; we aim to highlight the utility of this approach. Using our method, we provide substantive results regarding persistent COVID-19 symptoms in the HCT population during the early Omicron period of the SARS-CoV-2 pandemic. To accompany these results, we devise a sensitivity analysis approach investigating the degree to which the estimated time-to-symptom-resolution distribution changes under deviations from our key assumption. We also investigate baseline patient characteristics associated with time to resolution, that is, risk factors for delayed symptom resolution, using existing regression methods for current status data.
@@ Methods
In an ideal study with unlimited follow-up, we would observe for each participant the survey response time $Y_{*}$ — the time elapsed between positive SARS-CoV-2 test and survey response — and the event indicator $\Delta_{*}$ — a participant’s symptom status at the time of survey response — with $\Delta_{*}=1$ representing symptom resolution by time $Y_{*}$ and $\Delta_{*}=0$ otherwise. We denote by $c_{0}$ the investigator-determined time (since the positive test) at which follow-up ends for a participant, and consider as a nonrespondent any participant without a response by time $c_{0}$. The value of $c_{0}$ is the same for all participants in the study. We set $Y=\min(Y_{*},c_{0})$ as the smaller of the survey response time and $c_{0}$, and denote by $\Delta=\mathbbm{1}(Y_{*}<c_{0})\Delta_{*}$ the observed response indicator, which equals $1$ if a participant responded and reported resolution and 00 otherwise. The available data are considered to be $n$ independent and identically distributed observations $(W_{1},Y_{1},\Delta_{1}),(W_{2},Y_{2},\Delta_{2}),\ldots,(W_{n},Y_{n},\Delta_{n})$.
@@
Traditionally, under current status sampling, the survival function has been estimated based on the nonparametric maximum likelihood estimator (NPMLE), which can be computed using the pool adjacent violators algorithm used for isotonic (i.e., monotone) regression (Ayer et al.,, 1955). The key idea behind this approach is that estimation of the survival function can be reframed as a regression problem under a monotonicity constraint: the probability of symptom resolution by time $t$ is simply the proportion of individuals who have resolved symptoms among those with survey response time $t$, and this probability is necessarily nondecreasing as a function of survey response time. This estimator has the advantage of being distribution-free, in the sense that its validity does not rely on any assumption about the underlying shape of the true survival function, and of enabling the construction of valid confidence intervals derived from its large-sample properties (Groeneboom,, 1987). Nevertheless, an important limitation of the NPMLE approach is that it requires the survey response process to be entirely uninformative, in the sense that the response time itself does not carry information about the event time. This condition may fail to hold in practical settings, as there may exist baseline covariates — analogous to confounders in a causal inference setting — that inform the relationship between the event indicator and the survey response time. For example, participants with substantial comorbidities may be more likely to respond promptly to the survey and to have greater symptom persistence. We argue here that the framework of CIR introduced in Westling et al., (2020), which was developed to estimate a monotone causal dose-response curve, can be adapted to estimate the survival function with current status data under weaker conditions. Specifically, as opposed to the traditional NPMLE approach, this method only requires independence between the survey response and event times within strata defined by the recorded baseline covariates. Such independence condition is typically much more realistic, and thereby makes the analysis less likely to yield misleading conclusions due to assumption violations.
@@
are then obtained and regressed against $Y_{1},Y_{2},\ldots,Y_{n}$ over the domain $[t_{0},t_{1}]$ using isotonic regression. Technical details are provided in Section S1 of the Supplementary Material, including how it relates to the isotonization procedure of van der Vaart and van der Laan, (2006) in the absence of survey nonresponse. We denote the resulting estimator of $\psi(t)$ as $\psi_{n}(t)$. Use of these pseudo-outcomes rather than the simple event indicators accounts for possible dependence between $\Delta_{*}(t)$ and $Y_{*}$, as long as this dependence is explained entirely by $W$.
Roughly speaking, if at least one of $\mu$ and $g$ is estimated consistently, then $\psi_{n}(t)$ is itself consistent for $\psi(t)$, and if both $\mu$ and $g$ are estimated consistently at fast enough rates, $n^{\frac{1}{3}}\{\psi_{n}(t)-\psi(t)\}$ tends in distribution to a random variable with a standard Chernoff distribution (Groeneboom and Wellner,, 2001) scaled by a constant $\tau(t)$. This well-characterized large-sample distribution enables the construction of confidence intervals with correct coverage. The scale factor $\tau(t)$ of the limiting distribution, whose explicit form is given in Section S1 of the Supplementary Material, depends on $\mu$ and $g$, as well as on the derivative of the survival curve at $t$. To estimate $\tau(t)$, we require an estimate of this derivative, which can be obtained by applying a smoothing procedure to the estimated survival curve $\psi_{n}(t)$ and then using numerical differentiation. Once the scale factor has been estimated, with estimate denoted by $\tau_{n}(t)$, confidence intervals with nominal pointwise coverage $1-\alpha$ can be constructed as
$\left(\psi_{n}(t)-q_{1-\alpha/2}\tau_{n}(t)n^{-\frac{1}{3}},\psi_{n}(t)+q_{1-\alpha/2}\tau_{n}(t)n^{-\frac{1}{3}}\right),$
where $q_{\beta}$ denotes the $\beta$–quantile of the standard Chernoff distribution. A key strength of the CIR approach is that it allows for flexible estimation of $\mu$ and $g$ — using machine learning, for example — without sacrificing statistical inference. This reduces the risk of systematic bias due to the use of misspecified parametric models. In Section S4 of the Supplementary Material, we present results of a simulation study illustrating the relative insensitivity of our proposed procedure to the choice of nuisance estimators.
@@
We note that performing the CIR procedure using data from survey respondents alone — in other words, performing a complete case analysis — could result in bias, since whether a participant is considered a respondent is determined by their value of $Y_{*}$, which may depend on $W$ and thus be informative for $\Delta_{*}$. Symbolically, the estimand in a complete case analysis would be $1-E\{E(\Delta\,|\,W,Y=t)\,|\,Y<c_{0}\}$ rather than $1-E\{E(\Delta\,|\,W,Y=t)\}$. Therefore, fully removing nonrespondents — that is, participants with $Y_{*}\geq c_{0}$ — may induce selection bias, as the distribution of $W$ may differ among survey respondents and nonrespondents. To emphasize the difference between the complete case CIR and our proposed method, which includes all survey recipients, we refer to the latter as extended CIR.
Even with a rich set of baseline covariates $W$, there may be concern that conditional independence of $\Delta_{*}(t)$ and $Y_{*}$ is unlikely to hold. In Section S2 of the Supplementary Material, we detail a sensitivity analysis procedure designed to assess how deviations from conditional independence affect the estimates of $\psi(t)$ produced by the extended CIR procedure. The sensitivity analysis is based on an assumed copula model that encodes dependence between $\Delta_{*}(t)$ and $Y_{*}$ within covariate strata (Genest,, 1987). By varying the dependence parameter of the copula model, an investigator can observe how sensitive the CIR estimates are to violations of conditional independence. We note that the dependence parameter cannot be estimated from the observed data and must be specified by the investigator in order to estimate $\psi(t)$ under dependence within covariate strata.
@@
It is also of interest to understand the association between baseline covariates $W$ and symptom duration $T$ in order to identify potential predictors of delayed symptom resolution. In Section S3 of the Supplementary Material, we review regression models that have been adapted for use with current status data and discuss the difficulties that arise in making inference using these models. Focusing on the Cox model, we conduct a simulation study to evaluate the performance of the nonparametric bootstrap for making inference, finding that Wald-type bootstrap confidence intervals based on a normal approximation demonstrate good performance in modest sample sizes. The Cox model requires no special adaptation to account for nonresponse, as long as the regression model is correctly specified and includes all covariates that inform the response mechanism (Little and Rubin,, 2019).
@@
The HCT study was conducted at UW, a large public university in Washington state with a population composed of approximately 60,000 students and 32,000 faculty and staff. Beginning in autumn 2020, the HCT program provided free, voluntary PCR testing for SARS-CoV-2 and symptom surveillance for university affiliates. Individuals were sent daily messages via text or email to ask about new symptoms, exposures, and high-risk behaviors, any of which would trigger an invitation for PCR testing. Study participants were also able to walk-in for testing at any time, and for any reason, without an invitation. Once antigen testing became widely available in March $2022$, individuals were asked on their daily messages to report any antigen test results. The inclusion criteria for HCT eligibility are described in Section S5 of the Supplementary Material. Informed consent was signed electronically at the time of enrollment.
For this analysis, we used data collected from December $2021$ to June $2023$. During this time period, participants who had a positive PCR test or who reported a positive rapid antigen test for SARS-CoV-2 were invited to participate in a follow-up survey to assess for ongoing symptoms of COVID-19 a minimum of four weeks after the positive test.
@@
    Long COVID survey: Starting in December $2021$, all participants who tested positive for SARS-CoV-2 by PCR testing or those who attested to having had a positive rapid antigen test were contacted a minimum of $28$ days after their first positive test date and invited to participate in the long COVID survey. Each participant was allowed to fill out the survey only once per academic year. For participants with multiple responses, only the first is included in this analysis. The survey included the question “Are you still experiencing symptoms related to your COVID-19 illness (e.g., brain fog, fatigue, headaches, etc.)?” Each participant was free to choose when to complete the survey. Data collection was terminated on June $23$, $2023$.
@@
We implemented the extended CIR method using flexible machine learning tools to estimate $\mu$ and $g$. To estimate $\mu$, we used the Super Learner (van der Laan et al.,, 2007; Rose,, 2013), a form of stacked regression, with the optimal linear combination of learners selected to minimize $10$-fold cross-validated mean squared error — see Section S5 of the Supplementary Material for the algorithm library. To estimate $\pi$, we used the haldensify software package (Hejazi et al.,, 2022) with $10$-fold cross-validation used to select optimal tuning parameters. The derivative of the survival curve, which is required for confidence interval construction, was estimated by numerical differentiation of a monotone Hermite spline approximation of the estimated survival curve. Confidence intervals with nominal $95$% pointwise coverage were constructed using the 97.5th quantile of the Chernoff distribution.
@@
We aimed to estimate the survival function for symptom resolution between $t_{0}=30$ and $t_{1}=90$ days after the positive test. We therefore chose the maximum follow-up time to be $c_{0}=120$; participants with positive tests after February $24$, $2023$ were excluded to ensure all individuals in the analysis cohort had at least $120$ days of follow-up. On one hand, large values of $c_{0}$ result in more individuals excluded from the study period but make $t_{1}$ further away from the largest possible observed survey response times, where the statistical performance of isotonic regression and related procedures is known to degrade. On the other hand, small values of $c_{0}$ result in fewer individuals excluded from the study period but make $t_{1}$ closer to the largest possible observed survey response times. Our choice of $c_{0}$ was meant to achieve a balance between these competing considerations, and resulted in $t_{1}$ lying at approximately the $97.5$th percentile of observed survey response times; such choice performed well in a simulation study — see Section S4 of the Supplementary Material for details.
We also performed a sensitivity analysis using the procedure detailed in Section S2 of the Supplementary Material. We varied the copula association parameter to encode different degrees of conditional dependence between $T$ and $Y_{*}$ given $W$, corresponding to values of Kendall’s rank correlation coefficient ranging from -0.25 to 0.25. We used the same estimated nuisance functions as in the main analysis so that any observed variation in results can be directly ascribed to deviations from the conditionally uninformative response time assumption.
@@
To investigate the association between baseline risk factors and time from infection to symptom resolution, we fit a Cox proportional hazards regression model accounting for current status sampling using the icenReg software package (Anderson-Bergman,, 2017). Main terms were included in the model for each baseline covariate described above. Standard error estimates were calculated using a nonparametric bootstrap with $1000$ bootstrap replicates, and used to construct $95$% Wald-type confidence intervals and perform two-sided tests of the null regression coefficient hypothesis at a $0.05$ level. For categorical risk factors, multivariate Wald tests were used to test the joint null hypothesis that all of the corresponding regression coefficients were zero.